OpenStudy (anonymous):

Determine if the function f(x)=2x|x^3| is even, odd, or neither.

hartnn (hartnn):

a function is even when f(x) = f(-x) odd when f(x) = -f(-x) to get f(-x) just replace x in f(x) by -x !

hartnn (hartnn):

what you for as f(-x) for your f(x) ? could u find it ?

hero (hero):

Just replace x with -x. If you get the same expression afterwards, then the function is even. If not, then it is not even.

OpenStudy (anonymous):

Thanks! :D

hero (hero):

@Anacia23, you fully understand what to do here?

OpenStudy (anonymous):

@Hero Yeah. Just couldn't find my math notes..

hero (hero):

I see. Well, as long as you know what to do. So what did you conclude? Is the function even, odd, or neither?

OpenStudy (anonymous):

I got an odd function.

OpenStudy (zarkon):

\[f(-x)=f(x)=2(-x)|(-x)^3|\] \[=-2x|-(x)^3|\] \[=-2x|x^3|=-f(x)\]

OpenStudy (zarkon):

ignore the f(x) in the first line

OpenStudy (zarkon):

\[f(-x)=2(-x)|(-x)^3|\]

hartnn (hartnn):

yeah, f(x) is indeed odd function :)

OpenStudy (anonymous):

Thanks so much guys! And thanks @hartnn I thought so too!

hartnn (hartnn):

welcome ^_^